OpenStudy (anonymous):

WILL MEDAL AND FAN!! PLEASE HELP????? The revenue R (in dollars) from renting x apartments can be modeled by R = 2x(700 + 34x − x^2). (a) Find the additional revenue when the number of rentals is increased from 14 to 15. $ _________ (b) Find the marginal revenue when x = 14. $ _________ (c) Compare the results of parts (a) and (b). The answers are close. The answers are not close.

OpenStudy (kirbykirby):

is that \(2 \times (700 + 34x -x^2)\) , i.e 2 times ( ... ) or \(2x(700 +34x-x^2)\), i.e. 2x times ( ...)

OpenStudy (anonymous):

it's 2x not 2 times

OpenStudy (kirbykirby):

for a), you can find the revenue for x=14, and for x=15, and subtract off the two revenues. \(R_{14} = 2(14)(700 + 34(14) − (14)^2)\) \(R_{15} = 2(15)(700 + 34(15) − (15)^2)\) Then find the value \(R_{15}-R_{14}\)

OpenStudy (anonymous):

is it 2110?

OpenStudy (kirbykirby):

yes

OpenStudy (kirbykirby):

Marginal revenue then for b) is finding \(R\,'_{14}\), so the derivative evaluated at x = 14 \[\large R\,' _{14} =\left.\frac{d}{dx}2x(700+34x-x^2)\right|_{x=14}=\left.\frac{d}{dx}(1400x+68x^2-2x^3)\right|_{x=14}\\ =\large \left. 1400+68(2)x-2(3)x^2\right|_{x=14}=1400+68(2)(14)-2(3)(14)^2\]

OpenStudy (anonymous):

so part b is 2120?

OpenStudy (anonymous):

so part c would be close but not equal?

OpenStudy (kirbykirby):

i get 2128

OpenStudy (kirbykirby):

i suppose they would be considered close

OpenStudy (anonymous):

aww.. well they are still close and it really helped you walking me through this Thank you!

OpenStudy (kirbykirby):

:) yw